Calculate the molar solubility of Pb (IO3)2 in 1.0 x 10–4 M Pb (NO3)2. Ksp =2.5 x10-13
1
Expert's answer
2020-09-21T06:32:10-0400

In the solution of 1.0x10-4 M Pb(NO3)2 the concentrations of ions in solution is thus;

[Pb2+] = 1.0 x 10-4M

[NO3-] = 2x 1.0 x10-4 = 2.0 x 10-4 M


On addition of Pb(IO3)2, the solution will dissociate partly as;


Pb(IO3)2 "\to" Pb2+(aq) + 2IO3- (aq)


I = 1.0x10-4 0


C = X 2X

E = (1.0x10-4 + X ) 2X



Therefore; Ksp "=" "[Pb^2+]^1 2[IO3^-]^2"


"2.5x10^-13 = [1.0x10^-4 + X]^1[2X]^2"


The X in the [1.0x10-4 + X] is so small that it can be neglected/ignored because from its very low Ksp, it is a greatly insoluble salt


Thus our equation becomes;


"2.5x10^-13 = [1.0x10^-4]^1[2X]^2"


Since, [2X]2 = 4X2


and (1.0x10-4 )(4X2) = 4.0x10-4X2


Therefore, 2.5x10-13 = 4.0 x10-4X2


Dividing both sides by 4.0 x10-4 gives,


"\dfrac{2.5x10^-13}{4.0x10^-4}=\dfrac{4.0x10^-4X^2}{4.0x10^-4}"


6.25x10-10 = X2

Removing the exponent


(6.25x10-10)"^\dfrac{1}{2}" = X2"^\dfrac{1}{2}"


2.5x10-5 = X


Therefore the molar solubility of P(IO3)2 in 1.0 x 10–4 M Pb (NO3)2 is 2.5x10-5 M